a rectangular sheet of paper is 20.5 cm wide. when it is folded in half, with the crease running parallel to the 20.5 cm sides, the resulting rectangle is the same shape as the unfolded sheet. Find the length of the sheet, to the nearest tenth of a cm. (in Europe, the shape of notebook paper is determined by this similarity property.)
Answer by oscargut(2103)

x=length of the original paper
20.5=wide of the original paper
20.5=length of the 2nd paper
x/2=wide of the 2nd paper
the proportion between the sides are the same so
20.5/x = (x/2)/20.5
20.5/x=x/41
(20.5)41=x^2
x=sqrt(840.5)=29 (aprox)